OpenStudy (anonymous):

guys can you answer this pls? thankyou x^-1/2 -1/x^1/2 all over 3x^-1/3 -3/x^1/3 . . . thankyou .muchas gracias

OpenStudy (anonymous):

\[\frac{ x ^{\frac{ -1 }{ 2 } }-\frac{ 1 }{ x ^{\frac{ 1 }{ 2 }}} }{ 3x ^{-\frac{ 1 }{ 3 }}-\frac{ 3 }{ x ^{\frac{ 1 }{ 3 }} } }\] is this the sum?

OpenStudy (anonymous):

can u you use () to separate the value.am not sure if i have the right equation

OpenStudy (anonymous):

@xtel is this the right equation?

OpenStudy (anonymous):

@Razzputin yes ^^ thats it

OpenStudy (anonymous):

k kwl do you know how to start finding the answer to this?

OpenStudy (anonymous):

@reeshy2all the equation above that razzputin has replied. ^^

OpenStudy (anonymous):

@Razzputin i think add first the then find the lcd above .im not really sure

OpenStudy (anonymous):

Lol....Can u write..ur Question in Latex..)

OpenStudy (anonymous):

@Yahoo! the equation above that razzputin has replied

OpenStudy (anonymous):

can you not see the equation i posted @Yahoo! ? @xtel if you look at the numeratot do you know how to simplify a term that is \[\frac{ 1 }{ x }\]?

OpenStudy (anonymous):

\[x ^{-1/2} = \frac{ 1 }{ x ^{1/2} }\]

OpenStudy (anonymous):

Lol...i just Gave her a Hint :

OpenStudy (anonymous):

hope...nw u can solve @xtel

OpenStudy (anonymous):

IS THE CHAPTER ON FACTORIZATION

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{x} } - \frac{ 1 }{ \sqrt{x}}\]

OpenStudy (anonymous):

amp LOL. is the LCd x^-1/2 and -1/x^1/2 is x^1/2?

OpenStudy (anonymous):

THE SOLUTION IS O/NUMBER ,WHAT TOPIC IS THE CHAPTER ON.ALSO, DO U HAVE A RANGE FOR X

OpenStudy (anonymous):

no xtel if you are wanting to simplify it you need to make the exponet negative if you have 1/something like this: \[\frac{ 1 }{ x } = x ^{-1}\]

OpenStudy (anonymous):

@reeshy2all @Razzputin @Yahoo! Thankyou :))

OpenStudy (anonymous):

lol do you still need help?

OpenStudy (anonymous):

im just wndering if the answer will be 1 or 0 because ive got \[1/x _{2}^{1} - 1/x _{2}^{1} \div 3/x _{3}^{1} - 3/x _{3}^{1} \]

OpenStudy (anonymous):

@Razzputin just wonderin if it will become 0 lol